A dog running in an open field has components of velocity vx = 1.8 m/s and vy = -1.6 m/s at time t1 = 10.0 s . For the time interval from t1 = 10.0 s to t2 = 21.9 s , the average acceleration of the dog has magnitude 0.54 m/s2 and direction 30.0 ∘ measured from the +x−axis toward the +y−axis.

a)At time t2 = 21.9 s, what is the x-component of the dog's velocity?

b) At time t2 = 21.9 s , what is the y-component of the dog's velocity?

c)What is the magnitude of the dog's velocity?

d)What is the direction of the dog's velocity (measured from the +x−axis toward the +y−axis)?
Answer #1

a)

vx = 1.8 + (0.54 cos 30)* (21.9 - 10)

vx = 7.365 m/s

=====

b)

vy = - 1.6 + 0.54 sin 30 * (21.9 - 10)

vy = 1.613 m/s

=====

c)

v^2 =vx^2 + vy^2

v = 7.54 m/s

=====

d)

phi = arctan ( vy/vx)

phi = 12.353 anti clockwise from +x axis

=====

Comment in case any doubt, will reply for sure . Goodluck